# Uncertainty Quantification of Surrogate Models using Conformal Prediction

### v2

@@ Abstract
Data-driven surrogate models have shown immense potential as quick, inexpensive approximations to complex numerical and experimental modelling tasks. However, most surrogate models of physical systems do not quantify their uncertainty, rendering their predictions unreliable, requiring further validation. Though Bayesian approximations offer some solace in estimating the error associated with these models, they cannot provide guarantees, and the quality of their inferences depends on the availability of prior information and good approximations to posteriors for complex problems. This is particularly pertinent to multi-variable or spatio-temporal problems. Our work constructs and formalises a conformal prediction framework that satisfies marginal coverage for spatio-temporal predictions in a model-agnostic manner, requiring near-zero computational costs. We provide an extensive empirical study of the application of the framework to ascertain valid error bars that provide guaranteed coverage across the surrogate model’s domain of operation. The application scope of our work extends across a large range of spatio-temporal models, from solving partial differential equations to weather forecasting. Through the applications, the paper looks at providing statistically valid error bars for deterministic models, as well as crafting guarantees to the error bars of probabilistic models. Our conformal prediction formalisation provides guaranteed coverage of the surrogate model, regardless of model architecture, and its training regime and is unbothered by the curse of dimensionality.
@@ 1 Introduction
Partial Differential Equations (PDEs) governing physical processes are solved using complex numerical modelling simulation codes. Though these codes offer a mathematically-rigorous method of solving the relevant equations, the solutions are often limited to discretised domains and require computationally-expensive iterative solvers like finite-volume schemes and finite-element methods. Such simulation codes have become central to multiple scientific disciplines in biology (Hospital, Adam et al., 2015), engineering (Giudicelli et al., 2024), and climate science (Danabasoglu et al., 2020) (Lavin et al., 2021), but they are difficult to deploy for rapid, iterative modelling, which is often required while exploring a design space of solutions. Machine learning offers an alternative data-driven route for obtaining quick, inexpensive approximations to numerical simulations (Bertone et al., 2019; Karniadakis et al., 2021). Data-driven surrogate models are often built by distilling the spatio-temporal information characterised by the simulations into parameterised machine learning models. Due to the efficiency, cost-effectiveness, and relative accuracy of modelling, neural networks have become ubiquitous within scientific modelling, and they are of primary importance in tackling large-scale PDEs in climate (Lam et al., 2023; Kurth et al., 2023), computational fluid dynamics (Jiang et al., 2020; Pfaff et al., 2021), and nuclear fusion (van de Plassche et al., 2020; Gopakumar and Samaddar, 2020).
@@
As effective as these surrogate models (emulators) are, they remain approximations of the true physical system under study. There are often many layers of approximations as the simulation codes approximate the PDEs and the PDEs themselves are approximations to the true physical system. They often fail to quantify the uncertainty associated with their approximation, i.e. the approximation error of the emulator with that of the numerical code. Irrespective of the domain of operation they were trained on, models often claim confident outputs. This is problematic for two reasons: (a) without an assessment of the confidence of the output, erroneous predictions can lead to severe consequences downstream; (b) The cost of training these models can be prohibitively high, and without appropriate uncertainty quantification the overconfident predictions have limited utility in practice. Several works have attempted to provide uncertainty estimation for these surrogate models (Geneva and Zabaras, 2020; Alhajeri et al., 2022; Zou et al., 2024; Psaros et al., 2023), but they fail to provide statistical guarantees over the error bars attached to the prediction and to scale to complex scenarios (Abdar et al., 2021). They also require ensemble training Lakshminarayanan et al. (2017), extensive sampling (MacKay, 1992), or architectural modifications (Gal and Ghahramani, 2016), successively increasing the amount of computational resources required. There remains a concerning question of validating the surrogate model’s outputs for a specific downstream application.
@@
Conformal prediction (CP) (Vovk et al., 2005) provides a framework that allows us to compute statistically guaranteed error bars over pre-trained and fine-tuned models, i.e. the error bars are calibrated to provide the required coverage. Conformal prediction relies on the calibration of model performance across a specific dataset representative of the desired prediction distribution, and utilising the calibration measures to help provide valid error bars for the model output.
@@
In this paper, we conduct a thorough empirical study on conformal prediction for neural-network-based surrogate models to equip their predictions with calibrated, guaranteed error bars, even on out-of-training-distribution scenarios. We demonstrate guaranteed coverage across spatio-temporal domains (high-dimensional outputs), irrespective of the choice of model, data and training conditions. Our work provides a rigorous method to check the usefulness of a pre-trained surrogate model, its validity and applicability in an inference/production scenario.
We conduct an empirical study that demonstrates at scale the estimation of statistically guaranteed error bars for pre-trained and fine-tuned surrogate models using conformal prediction. Our focus lies on extending the conformal prediction framework to work for models predicting over a spatio-temporal domain. Through experimentation of increasing complexity, we demonstrate how we can provide guaranteed error bars for various kinds of neural-network-based surrogate models. Our experiments demonstrate that this method allows us to obtain guaranteed error bars with meaningful estimation even when we predict on data representing a different physical setting from that with which the model was trained on. Our study shows that irrespective of the dimensions of the output (our experiments explore models with output dimensions upwards of 20 million), we can get guaranteed coverage. We explore various methods of conformal prediction, comparing cost, performance, and associated architectural modifications. We conclude by stating that using conformal prediction we can obtain guaranteed marginal coverage across the spatio-temporal domain of interest, irrespective of the choice of model (both deterministic and probabilistic), training data and conditions.
Pathway to Addressing Climate Change
Machine-learning-based surrogate modelling is key to accelerating scientific simulation as it promises computational efficiency and allows for data-driven discovery at scale. Especially when modelling complex systems, such as computational fluid dynamics, nuclear fusion, and weather forecasting, computational speed of scientific simulation is important; furthermore, it is increasingly crucial to provide accurate, robust, and timely uncertainty estimates. In safety-critical systems Knight (2002), it is impervious that we supplement model predictions with calibrated uncertainty estimations. This can lead to improved decision-making processes in downstream tasks. The CP framework demonstrated in this paper pushes the envelope forward in uncertainty quantification for complex scientific modelling at scale, demonstrating industry-level safety-critical applications.
Solving differential equations is a crucial tool for being able to model dynamic systems. Neural-PDE solvers offer a quick, relatively inexpensive method to solve PDEs(Yin et al., 2023), allowing us to further understand the physical system and find optimum design points(Li et al., 2023a; Shukla et al., 2024). These Neural-PDE based surrogate models offer a new pathway to solve, understand and model physical systems. But as they become more ubiquitous across various application domains, ability to verify the accuracy of their predictions becomes pressingly important.
Nuclear fusion offers a solution to low-carbon, safe, renewable, and cheap energy production. Therefore, solving nuclear fusion at a commercial scale is a key component in decarbonising hard-to-abate industries (Lerede et al., 2023). However, several challenges remain, ranging from confinement(Degrave et al., 2022), and materials (Linke et al., 2019) to mitigating disruptions (Kates-Harbeck et al., 2019). Operating on an accelerated timeline and exponential costs, iterative test-based design becomes an infeasible option. Researchers are left to solve these challenges in-silico using simulations. However, within the context of fusion modelling, at the exascale, simulation costs grow exponentially and need surrogate models capable of quicker, cheaper approximations. Neural-PDE solvers have enabled this, allowing us to solve complex multi-physical systems several orders of magnitude faster than traditional numerical solvers (Pamela et al., 2024; Carey et al., 2024). As further developments are done on these models, it is pertinent to provide rigorous methods of quantifying the uncertainty associated with these models.
Climate change has led to the observation of extreme weather events across the globe (Ebi et al., 2021). Being able to forecast weather rapidly and cheaply allows us to take proactive rather than reactive measures to these climate change-induced events. Neural weather models learnt from both simulation and observed data have emerged as cheaper and faster alternatives to traditional numerical models. However, they suffer from the limitations found within the training data, such as sparse and inaccurate observations as well as the error growth in the numerical models (Sheshadri et al., 2021). Modelling the uncertainty associated with these neural weather predictions is crucial in an operational context (Bouallègue et al., 2024). Through this work, we propose a model-agnostic method that gives calibrated error bounds for all variables, lead times and spatial locations, while not requiring any modifications to the model albeit with negligible computational costs.
@@ 2 Conformal Prediction
Conformal prediction (Vovk et al., 2005; Shafer and Vovk, 2008) answers the following question: Given some arbitrary dataset $(X_{1},Y_{1}),(X_{2},Y_{2}),...,(X_{n},Y_{n})$, and some machine learning model $\hat{f}:\mathcal{X}\to\mathcal{Y}$ trained on this dataset, what is the accuracy of $\hat{f}$ at predicting the next true label $Y_{n+1}$ at query point $X_{n+1}$. CP (Vovk et al., 2005) extends the point prediction $\hat{y}$ of $\hat{f}$ to a prediction set $\mathbb{C}^{\alpha}$, which is guaranteed to contain the true label $Y_{n+1}$ with probability
@@
CP quantifies errors in machine learning since the inequality in Equation 1 is guaranteed regardless of the selected machine learning model and the training dataset $\left\{(X_{i},Y_{i})\right\}_{i=1}^{n}$, other than that the samples used for the calibration (estimating model performance) are exchangeable (a weaker form of i.i.d.).
There have been several variants of CP since its original proposal by (Vovk et al., 2005). Inductive conformal prediction (Papadopoulos, 2008), which we pursue in this work, splits the data into a training set (the usual training set for the underlying ML model) and a calibration set, used to construct prediction set $\mathbb{C}^{\alpha}$ for each $\alpha$ value. The prediction sets are constructed, such that Equation 1 holds, by comparing the trained ML model to the unseen calibration data using a nonconformity score.
@@
Figure 1: Inductive CP framework over a deterministic model (see AER in Section 2.2): (1) Perform calibration using a nonconformity metric (L1 error residual (AER) with $\hat{s}$ representing the calibration scores, $y_{c},\tilde{y}_{c}$ the calibration targets and predictions respectively). (2) Estimate the quantile corresponding to the desired coverage from the CDF of the nonconformity scores ($n$ represents the calibration sample size, $(1-\alpha)$ the desired coverage, $F_{\hat{s}}^{-1}$ the quantile function applied over the inverse CDF of nonconformity scores, $\hat{q}$ the quantile matching the desired coverage). (3) Apply the quantile to the model predictions to estimate the prediction sets ($\tilde{y}_{p}$, the model predictions and $\hat{q}$ the upper and lower bars for the predictions).
The inductive CP framework explored further in this paper follows a three-step procedure as outlined in Figure 1. Initially, we perform a calibration of the model outputs, where a nonconformity metric of the model’s performance is evaluated. This is followed by the quantile estimation, which is performed by obtaining the required quantile from the cumulative distribution of the nonconformity scores obtained in the previous calibration step. The final step involves estimating the prediction sets by applying the estimated quantile over the model predictions in the interested prediction regime.
2.1 Conformal Prediction over a Spatio-Temporal Domain
Though conformal prediction as a method of performing uncertainty quantification was introduced for fixed single-point outputs in (Vovk et al., 2005), it has been gaining popularity for usage across spatio-temporal data in the recent past (Sun, 2022). In (Stankeviciute et al., 2021), inductive conformal prediction is extended further to provide uncertainty guarantees on time-series data for RNNs, where (Xu and Xie, 2021) explores the same for dynamic time-series forecasting. In (Xu et al., 2023), the authors explore the CP formulation to model the evolution of wildfires. The CP framework is further extended to the operator space in (Ma et al., 2024), where the authors using quantile regression learn the operator mapping to learn the upper and lower error bars of the functional space and then provide guarantees to it by using the CP framework.
2.1.1 Overview
We consider models that describe the evolution of spatio-temporal field variables defined by physical processes, such as those described in numerical PDE modelling, nuclear fusion diagnostics, and weather forecasting. Each modelling task is perceived as an initial value problem, where the calibration and prediction sets are pairs of inputs and outputs characterised by the initial condition and their respective solutions. Within the case of PDE modelling, each datapoint is a single numerical simulation, whereas for experimental modelling (such as nuclear fusion and weather), the datapoint is a single forecast driven by an initial condition.
Figure 2: Cell-wise calibration of uncertainty using CP. We explore the case in Section 3.3.1, where we deploy a U-Net trained to model the wave equation in an out-of-training-distribution setting. The top row shows the ground truth, while the second row represents the prediction made by the surrogate model. The third row shows the 95% coverage (twice the standard deviation: error $\sim 2\sigma$) obtained by performing MC dropout, before performing any calibration (uncalibrated error). The uncertainty obtained post-calibration using conformal prediction is shown in the last row reflecting 95% coverage ($\alpha=0.05$). The calibration procedure is performed cell-wise, where the error bar associated with each cell is calibrated to provide the required coverage for that cell. The uncertainties obtained from MC Dropout are unrealistically small, while the calibrated error bars provide coverage by design. Therefore, our framework helps accurately identify the regions with high levels of uncertainty, targeting the regions of higher-order dynamics.
Within our framework, we perform calibration for each cell individually, estimating the marginal coverage for each cell of the spatio-temporal tensor as demonstrated in Figure 2, i.e., the coverage is guaranteed for each cell taken individually. CP is performed across each spatio-temporal point output by the model, resulting in upper and lower coverage bands for each point. Upon calibrating for the error bars cell-wise, we averaged over each to estimate the coverage across the simulation domain. Within the estimation of the prediction sets, we don’t consider the influence of adjacent field points and implicitly expect the model to extract that within the learning process. We expect the discretised spatio-temporal domain to be the same across the calibration and prediction set.
In this paper, we propose a formulation of the CP framework that provides statistically valid, dimensionality independent marginal coverage over spatio-temporal prediction models. We demonstrate the efficacy of our approach through comprehensive and rigorous experimentation across a range of spatio-temporal models, including complex multi-physics models and state-of-the-art neural weather and nuclear fusion models.
2.1.2 Definition
We define a parametric model $\tilde{Y}=\hat{f}(X)$ that learns to map the evolution of an initial temporal sequence of spatial fields ($X\in\mathbb{R}^{T_{\text{in}}\times N_{x}\times N_{y}\times N_{var}}$) to a later temporal sequence of spatial fields ($\tilde{Y}\in\mathbb{R}^{T_{\text{out}}\times N_{x}\times N_{y}\times N_{\text{var}}}$). The model inputs and outputs are characterised by 4D tensors, where $T_{\text{in}},T_{\text{out}}$ represents the temporal dimension of the initial states and forecast respectively, $N_{x}$ represents the $x$-dimension, $N_{y}$ represents the $y$-dimension and $N_{\text{var}}$ the field dimension (number of modelled variables). The calibration procedure is defined as $\hat{q}=\hat{C}(\tilde{Y},Y)$, utilising the model prediction ($Tilde{Y}$) and ground truth ($Y$) to estimate the quantile ($\hat{q}$) associated with the desired coverage. The operation is executed in a point-wise manner since both $\tilde{Y},Y,\hat{q}\in\mathbb{R}^{T_{\text{out}}\times N_{x}\times N_{y}\times N_{\text{var}}}$. The quantile is further utilised (as given in Section 2.2) to obtain the lower ($L$) and upper error bars ($U$) across each of the cells to form the prediction set $\mathbb{C}$, where $L$ and $U$ have the same dimensionality as $\hat{q}$. At evaluation for a prediction point $X_{n+1}$ with true label $Y_{n+1}$, the expectation of coverage, i.e. the prediction set, is guaranteed to satisfy

### v3

@@ Abstract
Data-driven surrogate models offer fast, inexpensive approximations to complex numerical and experimental systems but typically lack uncertainty quantification, limiting their reliability in safety-critical applications. While Bayesian methods provide uncertainty estimates, they offer no statistical guarantees and struggle with high-dimensional spatio-temporal problems due to computational costs and dependence on prior specification. We present a conformal prediction (CP) framework that provides statistically guaranteed marginal coverage for surrogate models in a model-agnostic manner with near-zero computational cost. Our approach handles high-dimensional spatio-temporal outputs by performing cell-wise calibration while preserving the tensorial structure of predictions. Through extensive empirical evaluation across diverse applications—including partial differential equations, magnetohydrodynamics, weather forecasting, and fusion diagnostics—we demonstrate that CP achieves empirical coverage with valid error bars regardless of model architecture (MLP, U-Net, FNO, ViT, GNN), training regime, or output dimensionality (spanning 32 to over 20 million dimensions). We evaluate three nonconformity scores (conformalised quantile regression, absolute error residual, and standard deviation) for both deterministic and probabilistic models, showing that guaranteed coverage holds even for out-of-distribution predictions where models are deployed on physics regimes different from their training data. Calibration requires only seconds to minutes on standard hardware, with prediction set construction incurring negligible computational overhead. The framework enables rigorous validation of pre-trained surrogate models for downstream applications without retraining, providing actionable uncertainty quantification for decision-making in scientific domains. While CP provides marginal rather than conditional coverage and assumes exchangeability between calibration and test data—limitations we demonstrate empirically through sensitivity analyses—our method circumvents the curse of dimensionality inherent in traditional uncertainty quantification approaches, offering a practical tool for the trustworthy deployment of machine learning in the physical sciences.
@@ 1 Introduction
Partial Differential Equations (PDEs) governing physical processes are solved using complex numerical simulation codes. While these codes offer mathematically rigorous solutions, they are limited to discretised domains and require computationally expensive iterative solvers such as finite-volume and finite-element methods. Such simulation codes are central to scientific disciplines in biology (Hospital_Adam2015-bw), engineering (giudicelli2024moose), and climate science (cesm2; simintelligence), but are difficult to deploy for rapid, iterative modelling required while exploring design spaces. Machine learning offers an alternative data-driven route for obtaining quick, inexpensive approximations to numerical simulations (Bertone2019; Karniadakis2021). Data-driven surrogate models distil spatio-temporal information from simulations into parameterised machine learning models. Due to their efficiency, cost-effectiveness, and relative accuracy, neural networks have become ubiquitous within scientific modelling, with primary importance in tackling large-scale PDEs in climate (lam2022graphcast; pathak2022fourcastnet), computational fluid dynamics (jiang2020meshfreeflownet; pfaff2021learning), and nuclear fusion (van_de_Plassche_2020; Gopakumar_2020).
@@
However, these surrogate models remain approximations of true physical systems, inheriting multiple layers of uncertainty from both the numerical codes and the PDE formulations themselves. Critically, they often fail to quantify their approximation error relative to the numerical code, producing overconfident outputs regardless of their training domain. This poses two key problems: (a) without confidence assessment, erroneous predictions can lead to severe downstream consequences; (b) high training costs are wasted if predictions lack actionable uncertainty quantification. While several works have attempted uncertainty estimation for surrogate models (GENEVA2020109056; ALHAJERI202234; zou2022neuraluq; Psaros2023), they fail to provide statistical guarantees over error bars and do not scale to complex scenarios (ABDAR2021243). Moreover, they require computationally expensive ensemble training lakshminarayanan2017simple, extensive sampling (bnns), or architectural modifications (gal2016dropout). Validating surrogate model outputs for specific downstream applications remains a pressing challenge.
@@
Conformal prediction (CP) (vovk2005algorithmic) provides a framework for computing statistically guaranteed error bars over pre-trained and fine-tuned models, i.e. error bars calibrated to provide required coverage. Conformal prediction relies on calibrating model performance across a dataset representative of the desired prediction distribution, then utilising these calibration measures to provide valid error bars for model outputs.
@@
In this paper, we conduct a thorough empirical study demonstrating that conformal prediction provides statistically guaranteed error bars for neural-network-based surrogate models across spatio-temporal domains, even in out-of-training-distribution scenarios. Through experiments of increasing complexity, we show guaranteed marginal coverage irrespective of model choice (deterministic or probabilistic), output dimensionality (up to 20 million dimensions), training data, or physical setting. We explore various conformal prediction methods, comparing their cost, performance, and architectural requirements. Our work provides a rigorous method to assess the usefulness, validity, and applicability of pre-trained and fine-tuned surrogate models for inference and production scenarios.
Applications to Critical Scientific Challenges
Machine-learning-based surrogate modelling accelerates scientific simulation through computational efficiency and enables data-driven discovery at scale. When modelling complex systems such as computational fluid dynamics, nuclear fusion and weather forecasting, both computational speed and accurate uncertainty estimates are crucial. In safety-critical systems safetycriticalsystems, supplementing model predictions with calibrated uncertainty estimations is imperative for improved decision-making in downstream tasks. The CP framework demonstrated here advances uncertainty quantification for complex scientific modelling at scale with industry-level safety-critical applications. Neural-PDE solvers offer quick, inexpensive PDE solutions(Yin_2023), enabling system understanding and optimal design point identification(li2023geometryinformed; SHUKLAdnoairfoil). As these models become ubiquitous, verifying prediction accuracy becomes pressingly important. Applications span nuclear fusion for low-carbon energy production (LEREDE2023101144; Degrave2022; PFC; Kates-Harbeck2019; pamela2024neuralpararealdynamicallytrainingneural; carey2024dataefficiencylongterm) to weather forecasting for proactive climate change response (cc_extremeweather; gcm_error_growth; ddw_risk). Through this work, we propose a model-agnostic method providing calibrated error bounds for all variables, lead times, and spatial locations, requiring no model modifications with negligible computational costs.
Outline
The paper is structured as follows: Section 2 introduces the inductive CP framework, spatio-temporal data, associated exchangeability assumptions, and the mathematical extension of CP to spatio-temporal domains. Section 3 presents experiments deploying our CP framework across diverse modelling tasks, from PDEs to climate modelling. Section 4 discusses the framework’s strengths and limitations, and Section 5 concludes.
@@ 2 Conformal Prediction
Conformal prediction (CP) (vovk2005algorithmic; shafer2008tutorial) addresses a fundamental question in machine learning: given a dataset $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})$ and a trained model $\hat{f}:\mathcal{X}\to\mathcal{Y}$, how accurate is $\hat{f}$ at predicting the true label $Y_{n+1}$ for a new query point $X_{n+1}$? CP extends point predictions $\hat{y}$ to prediction sets $\mathbb{C}^{\alpha}$ that contain the true label with guaranteed probability:
@@
This coverage guarantee holds regardless of the model architecture or training procedure, requiring only that calibration samples are exchangeable (a weaker form of i.i.d)(vovk2005algorithmic).
Several CP variants exist, with inductive conformal prediction (papadopoulos2008inductive) being particularly efficient for neural networks. This approach splits data into a training set (for model training) and a calibration set (for constructing prediction sets $\mathbb{C}^{\alpha}$). The prediction sets satisfy Equation 1 by comparing model outputs to calibration data using a nonconformity score, i.e. a metric quantifying prediction error.
@@
Figure 1: Inductive CP framework using Absolute Error Residual (AER) nonconformity scores (see Section 2.2): (1) Calibrate by computing nonconformity scores ($\hat{s}$) from calibration predictions ($\tilde{y}_{c}$) and targets ($y_{c}$). (2) Estimate the quantile ($\hat{q}$) for desired coverage $(1-\alpha)$ using $n$ calibration samples and the inverse CDF $F_{\hat{s}}^{-1}$. (3) Construct prediction sets by applying $\hat{q}$ to test predictions ($\tilde{y}_{p}$).
The inductive CP framework operates through three fundamental steps (Figure 1). First, nonconformity scores $s(x,y)$ are computed on calibration data, quantifying the disagreement between predictions and ground truth—larger scores indicate worse model performance. Second, the $(1-\alpha)$-quantile $\hat{q}$ of these scores is estimated, providing the threshold for constructing prediction sets. Third, for any test point $x_{\text{test}}$, the prediction set $\mathcal{C}(x_{\text{test}})=\{y:s(x_{\text{test}},y)\leq\hat{q}\}$ includes all outputs whose nonconformity scores fall below $\hat{q}$ (gentle_introduction_CP). Critically, while the coverage guarantee in Equation 1 holds universally, regardless of model quality, the usefulness of the prediction sets depends entirely on the choice of nonconformity score function. Well-designed scores that accurately rank prediction difficulty yield tight intervals for easy inputs and wider intervals for challenging ones; poorly chosen scores produce uninformative but still valid prediction sets.
2.1 Conformal Prediction for Spatio-Temporal Data
While originally developed for single-output predictions (vovk2005algorithmic), CP has been extended to spatio-temporal domains (sun2022conformal; conformaltimeserires; cp_dynamic_timeseries; CP_Wildfire; ma2024calibrated). We consider models predicting the evolution of spatio-temporal fields in physical systems as found in PDE modelling, weather forecasting, and fusion diagnostics. Each modelling task is formulated as an initial value problem where calibration and prediction sets consist of input-output pairs characterised by initial conditions and their solutions.
Figure 2: Cell-wise uncertainty calibration using CP for a U-Net modelling the wave equation in an out-of-distribution setting (Section 3.3.1). Rows show: ground truth, model prediction, uncalibrated 95% coverage from MC dropout ($2\sigma$), and calibrated 95% coverage ($\alpha=0.05$) from CP. Cell-wise calibration provides guaranteed coverage for each spatial location. MC dropout produces unrealistically small uncertainties, while CP error bars correctly identify regions of high uncertainty corresponding to complex dynamics.
Our framework performs calibration independently for each cell of the spatio-temporal tensor (Figure 2), providing marginal coverage guarantees at every spatial and temporal location. This cell-wise approach yields upper and lower bounds for each point without explicitly modelling spatial correlations, instead relying on the neural network to capture spatial dependencies during training. The discretised spatio-temporal grid must remain consistent between calibration and prediction sets. We formalise this approach to provide statistically valid, dimension-independent marginal coverage for high-dimensional outputs, as demonstrated across diverse applications, including multi-physics systems and operational weather models.
2.1.1 Mathematical Formulation
Consider a model $\hat{f}$ mapping initial temporal sequences of spatial fields $X\in\mathbb{R}^{T_{\text{in}}\times N_{x}\times N_{y}\times N_{\text{var}}}$ to future sequences $\tilde{Y}=\hat{f}(X)\in\mathbb{R}^{T_{\text{out}}\times N_{x}\times N_{y}\times N_{\text{var}}}$, where $T_{\text{in}},T_{\text{out}}$ denote input and output time steps, $N_{x},N_{y}$ are spatial dimensions, and $N_{\text{var}}$ is the number of variables. The calibration procedure $\hat{q}=\hat{C}(\tilde{Y},Y)$ uses model predictions $\tilde{Y}$ and ground truth $Y$ to compute quantiles $\hat{q}\in\mathbb{R}^{T_{\text{out}}\times N_{x}\times N_{y}\times N_{\text{var}}}$ in a point-wise manner. These quantiles define lower ($L$) and upper ($U$) bounds forming the prediction set $\mathbb{C}$, with $L$ and $U$ sharing the dimensionality of $\hat{q}$. For a test point $X_{n+1}$ with true label $Y_{n+1}$, the coverage guarantee becomes: